what is the vertex of the parabola x^2-4y+8=0
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

what is the vertex of the parabola x^2-4y+8=0
---
Solve for "y":
4y = x^2+8
---
y = (1/4)x^2 +2
--------
Vertex occurs where x = -b/2a = 0/(1/2) = 0
---
If x = 0, y = 2
---------------------
Vertex: (0,2)
=================
Cheers,
Stan H.
=================


Answer by Alan3354(69443)

what is the vertex of the parabola x^2-4y+8=0
----------
x^2-4y+8=0
y = (1/4)x^2 + 2
The vertex is on the axis of symmetry, which is x = -b/2a
x = 0
@ x = 0, y = 2
Vertex at (0,2)